# JAX-MPM: A Learning-Augmented Differentiable Meshfree Framework for GPU-Accelerated Lagrangian Simulation and Geophysical Inverse Modeling

### v1

@@ 3.3.1 Computational speedup using JIT and loop optimization
As summarized in Section 2.2, MPM simulations consist of three core computational stages repeated at each time step: particle-to-grid (P2G) transfer, grid-based updates, and grid-to-particle (G2P) transfer. These operations are inherently parallel over particles and grid nodes, making them ideal targets for optimization. JAX-MPM accelerates these stages using two key strategies: just-in-time (JIT) compilation for kernel fusion and vectorization to parallelize computation across simulation entities.
@@
Remark 3 (Vectorized parallelism). To fully exploit the parallel structure of particle and grid computations, JAX-MPM combines jax.vmap and vectorized array operations. Besides using vmap for function-level operations such as polar decomposition, we apply batched array operations—e.g., scatter_add for momentum transfer in P2G—to update all grid nodes and particles in parallel. This eliminates explicit loops and enables efficient execution on both CPUs and GPUs.
@@ 3.3.2 Memory optimization via checkpointing and segmentation
To enable memory-efficient training over long simulation horizons, JAX-MPM incorporates two complementary techniques to reduce memory usage during reverse-mode automatic differentiation: checkpointing of substeps and segmentation of the simulation loop.
Remark 4 (Checkpointing grid operations). Unlike fully Lagrangian methods such as position-based dynamics, MPM adopts a hybrid Eulerian–Lagrangian formulation, where particles interact with a background grid. This incurs high memory costs during backpropagation, as intermediate grid states must be stored. To reduce memory usage, we apply gradient checkpointing: intermediate buffers associated with grid states are discarded during the forward pass and recomputed during the backward pass, retaining only particle-level outputs (e.g., $\bm{x}_{p}^{t}$, $\bm{v}_{p}^{t}$). Similar strategies have been used in prior work [49, 101]. Within the JAX-MPM framework, we implement this using jax.checkpoint (also known as jax.remat), which re-materializes intermediate computations in substep during backpropagation. A representative implementation is shown below:
@@
v, x, C, F = carry
base, fx, w = pre_compute(x)
F, gv_in, gm_in = p2g(v, x, C, F, base, fx, w)
@@
v, x, C = g2p(gv_out, v, x, base, fx, w)
return (v, x, C, F), None
Remark 5 (Segmented simulation loop with jax.scan). To reduce memory and compilation overhead, the simulation loop is divided into $n$ segments, each containing $N_{t}/n$ time steps. Each segment is executed using jax.scan and wrapped with jax.checkpoint, allowing intermediate values within the segment to be discarded during the forward pass and recomputed during the backward pass. As a result, peak memory usage is determined by the buffer size needed to recompute a single segment, reducing memory complexity from $\mathcal{O}(N_{t})$ to $\mathcal{O}(N_{t}/n)$, where $N_{t}$ is the total number of time steps. An example implementation is shown below:
@jax.remat
@@
Figure 3: Backward propagation of JAX-MPM with $N_{t}$ time steps
4 Numerical results of large scale simulation
@@ 4.1 Validation of JAX-MPM for 2D cases
4.1.1 Simulation of 2D dam break problem
Figure 4: (a): Configuration of 2D shallow dam break; (b): Configuration of quasi-2D granular collapse.
Accurate simulation of free-surface dynamics is essential for modeling geophysical processes such as floods, dam breaks, and debris surges, where the interaction of inertia, gravity, and evolving boundaries governs flow behavior.
To evaluate the capability of JAX-MPM in capturing such phenomena, we simulate a classical two-dimensional dam break problem under shallow-water assumptions, following the setup in [52]. As illustrated in Fig. 4a, a rectangular column of water with initial height $H_{0}=100\,\mathrm{mm}$ and length $L_{0}=1000\,\mathrm{mm}$ is released at $t=0$ by removing a gate on the right boundary. The fluid is modeled as a weakly incompressible Newtonian fluid with zero viscosity and friction, allowing direct comparison with analytical results. The simulation domain is discretized with a grid size of $\Delta h=0.004\,\mathrm{m}$, and each cell contains 4 particles. The numerical sound speed is $c_{n}=35\,\mathrm{m/s}$ and $\Delta t=1\times 10^{-5}$ s is set to maintain stability in weakly incompressible flow. Slip boundary conditions are applied to all domain boundaries.
This benchmark admits an analytical solution, providing a rigorous test for validating transient flow front propagation and surface evolution. The horizontal position of the flow front at time $t$ is given by shallow-water theory [89, 52]:
@@ 4.1.1 Simulation of 2D dam break problem
We test three different transfer schemes, including FLIP, TPIC, and APIC, within the JAX-MPM framework (see Section 2.2.3). As shown in Fig. 5, all methods yield results in close agreement with the analytical shallow-water solution overall. Small discrepancies appear during the early stage ($t<0.3\,\mathrm{s}$), primarily due to finite-depth effects and vertical accelerations that violate the shallow-water assumptions. As the flow becomes shallower ($t\geq 0.3\,\mathrm{s}$), the numerical results align more closely with the analytical solution. These observations are consistent with prior findings in [52], confirming the correctness of our implementation.
Figure 5: Simulation of 2D shallow dam break problem by JAX-MPM with (a) FLIP, (b) TPIC and (c) APIC.
@@ 4.1.2 Simulation of quasi-2D granular collapse: aluminum-bar assemblage
Here, we consider a granular flow benchmark involving the collapse of an aluminum-bar assemblage, following the experimental setup of Bui et al. [15]. The material behavior is governed by the Drucker–Prager elastoplastic model, as described in Sections 2.3 and A. The material is non-cohesive ($c=0$ kPa), with density $\rho=2650\,\mathrm{kg/m}^{3}$, bulk modulus $K=0.7\,\mathrm{MPa}$, Poisson’s ratio $\nu=0.3$, internal friction angle $\varphi=19.8^{\circ}$, and dilatation angle $\psi=0$.
[…]
Fig. 6a shows the contour plot of the equivalent plastic strain $\varepsilon_{eq,p}$ with peak values concentrated near the base of the aluminum-bar assemblage. These high-strain regions propagate upward, forming a localization pattern consistent with typical shear band development observed in granular collapse scenarios (cf. Fig. 15 in [74]). Fig. 6b presents the predicted free surface and failure surface, defined by a threshold of $\varepsilon_{eq,p}>0.03$, both of which show good agreement with experimental observations in [15].
These results demonstrate that the Drucker–Prager model in JAX-MPM accurately captures the key features of granular flow behavior, including shear localization, material failure, and free-surface evolution. Moreover, this benchmark highlights the robustness of the framework in simulating elastoplastic deformation and large-strain dynamics in geomechanical materials.
Figure 6: Simulation results of the aluminum-bar assemblage collapse using JAX-MPM. (a) Contour plot of equivalent plastic strain $\varepsilon_{eq,p}$, illustrating shear band development. (b) Comparison with experimental data from Bui et al. [15], where the green dots and black triangles denote the experimentally observed failure surface and free surface, respectively. The simulated failure surface is identified by a threshold $\varepsilon_{eq,p}>0.03$.
@@ 4.2 Application of JAX-MPM for 3D cases
4.2.1 Simulation of 3D dam break problem
We simulate a three-dimensional (3D) dam break scenario to evaluate the ability of JAX-MPM to capture complex free-surface flow around embedded structures.
As illustrated in Fig. 7, the computational domain is defined as [0, 1.0]×[0, 0.5]×[0, $[0,\,1.0]\times[0,\,0.5]\times[0,\,0.4]\,\mathrm{m}^{3}$. The initial dimensions of the water column are $H_{0}=0.2\,\mathrm{m}$ (height), $L_{0}=0.3\,\mathrm{m}$ (length), and $W_{0}=0.5\,\mathrm{m}$ (width). The fluid is modeled as a weakly compressible Newtonian fluid with a dynamic viscosity of $\mu=1.01\times 10^{-3}\,\mathrm{Pa\cdot s}$ and a numerical sound speed of $c_{n}=35\,\mathrm{m/s}$. The two rigid square columns, introduced as internal obstacles, have side lengths of 0.04 m and 0.06 m, positioned at $(0.5,0.3,0.0)\,\mathrm{m}$ and $(0.6,0.1,0.0)\,\mathrm{m}$, respectively.
A frictionless contact model employing a predictor–corrector scheme [10, 47] is used to prevent fluid penetration into the stationary rigid structures. Given the rigidity of these obstacles, enforcing the contact simplifies to eliminating the normal component of the fluid velocity at the fluid-structure interface:
@@ 4.2.1 Simulation of 3D dam break problem
Figure 7: Configuration of 3D dam break simulation with rigid internal obstacles.
@@
Besides capturing physically realistic flow dynamics qualitatively, JAX-MPM showcases high computational efficiency. The simulation presented in Fig. 8, featuring over 1.9 million particles and a $200\times 100\times 80$ grid, completes 1000 time steps in approximately 8 seconds (see Table 1). This computational performance highlights the scalability of the JAX-based implementation, which benefits from just-in-time (JIT) compilation and GPU acceleration. These features collectively enable fast and stable execution of large-scale geotechnical simulations, positioning JAX-MPM as a powerful computational tool for practical and high-performance modeling.
Figure 8: 3D dam break simulation with with rigid internal obstacles.
Table 1: Summary of simulation configurations and computational costs. The label “sim” refers to forward simulations, with runtimes reported for both single precision (float32) and double precision (float64), formatted as “float32 / float64”. The label “sim + grad” denotes inverse simulations involving both forward and gradient computations, performed exclusively in double precision (float64). “Sec/1000 steps” indicates the wall time (in seconds) for every 1000 time steps.
Example Particle # Grid resolution Param # Sec/1000 steps $\Delta h$
Dam break (Fig. 8) 1,920,000 $200\times 100\times 80$ N/A 8.2/35.1 (sim) $1/200$
@@
To evaluate the capability of JAX-MPM in simulating large-scale granular flow behaviors and capturing the influence of geometric factors, we consider a three-dimensional granular column collapse scenario. The primary goal is to investigate how the initial geometry, characterized by different aspect ratios $a=H_{0}/L_{0}$ affects the dynamics, runout distance, and failure mechanisms of the granular material. Here, $H_{0}$ and $L_{0}$ denote the initial height and length of the granular column, respectively. A total of 200,000 time steps are executed, corresponding to 2 s of simulation time with a time step size of $\Delta t=10^{-5}\,\text{s}$. The material parameters remain the same as in Section 4.1.2.
In this test, we select four aspect ratios for systematic comparison: $a=0.5,1.0,2.0,$ and $3.0$. As shown in Fig. 9, higher aspect ratios result in a larger portion of the granular column experiencing free fall, thereby decreasing the size of the static basal region and increasing lateral spreading. The final runout distance $L_{f}$ is measured and normalized by the initial base length $L_{0}$, deinfed as $d_{n}=(L_{f}-L_{0})/L_{0}$. The normalized runout distance increases from approximately 2.05 at $a=0.5$ to 3.97 at $a=1.0$, 7.76 at $a=2.0$, and 10.65 at $a=3.0$. While runout increases with the aspect ratio, the growth rate diminishes at higher values, reflecting a transition from linear to sublinear scaling. This trend is consistent with experimental observations reported in [87], where the normalized runout distance scales approximately as $d_{n}\propto a$ for small $a$ and $d_{n}\propto a^{0.7}$ for larger $a$.
@@
To evaluate the computational performance of JAX-MPM, we benchmark the 3D granular column collapse simulation against the open source C++ solver CB-Geo [60]. Simulations are performed at five resolutions, with grid spacings ranging from $\Delta h=0.015$ to $\Delta h=0.004$, corresponding to particle counts from approximately $1.7\times 10^{4}$ to $9.0\times 10^{5}$. Wall-clock times are measured for every 1000 simulation steps across the following computational configurations: (1) CB-Geo on multiple CPUs and (2) JAX-MPM on GPU. CPU benchmarks are conducted on an AMD EPYC 7763 processor with 128 threads, whereas GPU benchmarks utilize an NVIDIA A100 accelerator.
It should be noted that CB-Geo employs a Mohr–Coulomb material model, as the built-in solver does not currently support Drucker–Prager plasticity. Although Mohr–Coulomb models are generally more computationally demanding due to their nonsmooth yield surfaces, their computational complexity remains comparable to that of Drucker–Prager models. Therefore, the presented benchmarking results still provide valuable insight into the relative computational scalability and efficiency of both solvers on CPU and GPU platforms.
As shown in Fig. 10, JAX-MPM executed on GPU substantially outperforms CB-Geo on CPU configurations across all tested resolutions, exhibiting the most favorable scaling behavior. At the highest resolution ($\Delta h=0.004$, 900,000 particles), GPU-accelerated JAX-MPM completes 1000 steps in roughly 9 seconds, representing a speedup more than 7$\times$ compared to CB-Geo running on CPU.
The strong GPU performance and scaling are enabled by JAX-MPM’s effective utilization of GPU acceleration, just-in-time (JIT) compilation, and fused kernel execution (see Section 3.3.1). These features significantly enhance the solver’s capability for moderate to large-scale geomechanical problems, highlighting the computational advantages of differentiable simulation frameworks built on modern array-programming infrastructures.
Figure 9: 3D granular column collapse with varying aspect ratios. 2D deposit profiles are shown at $t=0.4\,\text{s},\,0.8\,\text{s},\,1.2\,\text{s},\,\text{and}\,\text{final deposit}$ for initial aspect ratios $a=H_{0}/L_{0}=0.5,1.0,2.0,3.0$. Higher aspect ratios lead to greater vertical collapse and longer runout distances due to enhanced free-fall and mass mobilization.
@@
In this example, we investigate the collapse of a 3D granular cylinder with varying internal friction angles $\varphi=20^{\circ},25^{\circ},35^{\circ},$ and $40^{\circ}$. The simulation runs for 1 s using a time step size of $\Delta t=1\times 10^{-5}\,\text{s}$, totaling 100,000 time steps. The Drucker–Prager elastoplastic constitutive model is employed, with all other material parameters remaining identical to those detailed in Section 4.1.2. Friction significantly influences shear resistance, flow mobility, and the final deposit geometry in granular flows.
As shown in Fig. 11, increasing the friction angle results in steeper and more confined final deposits with shorter runouts, reflecting increased resistance to shear deformation. These results align well with theoretical predictions and prior experimental observations reported in the literature [59, 57].
The JAX-MPM simulation involves over 2.7 million particles on a $400\times 400\times 100$ computational grid (approximately 16 million grid nodes) and completes every 1000 simulation steps in approximately 22 seconds (single precision) and 98 seconds (double precision) using a single NVIDIA A100 GPU. For comparison, a recent high-performance Fortran-based MPM solver employing dynamic load balancing (DLB) reported a runtime of 688 seconds for a similar scenario, involving roughly the same number of particles and comparable material models, with a grid resolution of $280\times 280\times 120$ (approximately 9.4 million grid nodes), executed on 288 CPU cores [45]. Without DLB, the runtime increases significantly to approximately 13,803 seconds. Relative to the Fortran implementation with DLB, JAX-MPM achieves roughly a 7×7$\times$7 × speedup, and nearly 140×140$\times$140 × over the baseline version without DLB. These benchmarks underscore the exceptional computational efficiency and scalability of the JAX-MPM framework, particularly when leveraging modern GPU acceleration for large-scale geomechanical simulations.

### v2

@@ 3.3.1 Computational speedup using JIT and loop optimization
As summarized in Section 2.2, MPM simulations consist of three core computational stages repeated at each time step: particle-to-grid (P2G) transfer, grid-based updates, and grid-to-particle (G2P) transfer. These operations are inherently parallel over particles and grid nodes, making them ideal targets for optimization. JAX-MPM accelerates these stages using two key strategies: just-in-time (JIT) compilation to enable kernel fusion and vectorization to parallelize computation across simulation entities.
@@
Remark 2 (Vectorized parallelism). To exploit particle and grid parallelism, JAX-MPM combines jax.vmap with vectorized array primitives. For example, polar decomposition is applied across particles with vmap, while momentum transfer in P2G uses indexed scatter-add updates (x.at[idx].add(val)) over all affected grid nodes. This removes explicit Python loops and enables efficient execution on both CPUs and GPUs.
@@ 3.3.2 Memory optimization via checkpointing and segmentation
Reverse-mode AD incurs large memory costs by storing intermediates, especially in grid-based operations. JAX-MPM mitigates this using checkpointing and segmented backpropagation.
Figure 3: Backward propagation with checkpointed grid operations in JAX-MPM. During the forward pass, only particle states are stored while grid buffers are discarded. During backpropagation, jax.checkpoint recomputes the grid operations, reducing memory usage in reverse-mode automatic differentiation (jax.grad).
Remark 3 (Checkpointing grid operations). The particle–grid interactions, within MPM’s hybrid Eulerian-Lagrangian formulation, require substantial buffer storage during backpropagation. To reduce memory usage, we apply gradient checkpointing: intermediate grid states are discarded during the forward pass and recomputed during the backward pass, while only particle-level outputs (e.g., $\bm{x}_{p}^{t}$, $\bm{v}_{p}^{t}$) are retained. Similar strategies have been used in prior work [48, 100]. Within the JAX-MPM framework, we implement this using jax.checkpoint (also known as jax.remat), which rematerializes intermediate computations in substep during backpropagation, as illustrated in Fig. 3. A representative implementation is shown below:
@@
v, x, rho, v_grad = carry
base, fx, w, dw = pre_compute(x)
rho, gv_in, gm_in = p2g(v, x, rho, v_grad, base, fx, w)
@@
v, x, v_grad = g2p(gv_out, v, x, base, fx, w, dw)
return (v, x, rho, v_grad), None
Remark 4 (Segmented simulation loop with jax.scan). To further reduce peak memory, the simulation loop is partitioned into $n$ segments, each containing $N_{t}/n$ time steps. Each segment is executed using jax.scan and wrapped with jax.checkpoint, allowing intermediate values within the segment to be discarded during the forward pass and recomputed during the backward pass. As a result, peak memory usage is determined by the buffer size needed to recompute a single segment, reducing memory complexity from $\mathcal{O}(N_{t})$ to $\mathcal{O}(N_{t}/n)$, where $N_{t}$ is the total number of time steps and $N_{t}/n$ is the segment length. An example implementation is shown below:
@jax.checkpoint
@@
4 Numerical results of predictive simulation
@@ 4.1 Validation of JAX-MPM for 2D cases
4.1.1 Simulation of 2D dam-break problem
Figure 4: (a): Configuration of the 2D shallow dam-break problem [51]; (b): Schematics of the quasi-2D granular collapse simulation (corresponding to the experimental configuration of aluminum-bar assemblage [14]).
Accurate simulation of free-surface dynamics is essential for geophysical applications such as floods, dam breaks, and debris surges. To validate the capability of JAX-MPM in capturing such behavior, we simulate a classical two-dimensional dam-break problem under shallow-water assumptions (Fig. 4a), following the setup in [51]. The fluid is modeled as a weakly incompressible Newtonian fluid with zero viscosity and friction, enabling direct comparison with analytical results.
As illustrated in Fig. 4a, a rectangular column of water with initial height $H_{0}=100\,\mathrm{mm}$ and length $L_{0}=1000\,\mathrm{mm}$ is released at $t=0$ by removing a gate on the right boundary. The simulation domain is discretized with a grid size of $\Delta h=0.004\,\mathrm{m}$, and each cell contains 4 particles. The numerical sound speed is $c_{n}=35\,\mathrm{m/s}$ and $\Delta t=1\times 10^{-5}$ s is set to ensure numerical stability. Slip boundary conditions are applied to all domain boundaries.
This benchmark admits an analytical solution that characterizes the transient flow front propagation and surface evolution. According to shallow-water theory [88, 51], the horizontal position of the front at time $t$ is given by
@@ 4.1.1 Simulation of 2D dam-break problem
Three transfer schemes, including FLIP, TPIC, and APIC (see Section 2.2), are evaluated within the JAX-MPM framework for this benchmark. As shown in Fig. 5, all schemes produce results in close agreement with the analytical shallow-water solution. Minor discrepancies observed during the early stage ($t<0.3\,\mathrm{s}$) are primarily attributed to finite-depth effects and vertical accelerations that violate shallow-water assumptions. As the flow becomes shallower ($t\geq 0.3\,\mathrm{s}$), the numerical results exhibit improved agreement with theory. These observations are consistent with prior findings in [51], confirming the correctness of our implementation.
Figure 5: Simulation results of 2D shallow dam-break problem by JAX-MPM with (a) FLIP, (b) TPIC and (c) APIC.
@@ 4.1.2 Simulation of quasi-2D granular collapse: Aluminum-bar assemblage
As another 2D benchmark, we simulate the collapse of an aluminum-bar assemblage, following the experimental setup of Bui et al. [14]. The material behavior is governed by the Drucker–Prager elastoplastic model, as described in Section 2.3 and A. The material is assumed to be non-cohesive ($c=0$ kPa), with density $\rho=2650\,\mathrm{kg/m}^{3}$, bulk modulus $K=0.7\,\mathrm{MPa}$, Poisson’s ratio $\nu=0.3$, internal friction angle $\varphi=19.8^{\circ}$, and dilatation angle $\psi=0^{\circ}$.
As shown in Fig. 4b, the initial granular column has dimensions $L_{0}=0.2\,\text{m}$ and $H_{0}=0.1\,\text{m}$, and is initially confined by a vertical gate. To approximate a quasi-2D condition, the simulation is performed under a plane strain assumption within a computational domain $(x,y,z)\in[0,\,0.6]\,\times[0,\,0.002]\,\times[0,\,0.12]\,\mathrm{m}^{3}$, discretized into a uniform grid of $300\times 1\times 60$ cubic cells. In the JAX-MPM simulation, no-slip boundary conditions are applied at the base, while slip conditions are imposed on all other boundaries. The FLIP transfer scheme is employed with a fixed time step $\Delta t=1\times 10^{-5}$ s over a total simulated duration of $0.65\,\mathrm{s}$.
Figure 6a shows the contour of the simulated equivalent plastic strain $\varepsilon_{eq,p}$, with peak values concentrated near the base and propagating upward to form localized shear bands. This is consistent with typical shear localization behavior observed in granular collapse (cf. Fig. 15 in [73]). Fig. 6b presents the predicted free surface and failure surface, defined by a threshold of $\varepsilon_{eq,p}>0.03$, both showing good agreement with experimental observations in [14].
These results demonstrate that the proposed JAX-MPM framework with the Drucker–Prager model accurately captures key features of granular flow behavior, including shear localization and free-surface evolution. Moreover, this benchmark highlights the robustness of our framework in modeling elastoplastic deformation and large-deformation dynamics in geological materials.
Figure 6: Simulation results of the aluminum-bar assemblage collapse using JAX-MPM. (a) Contour plot of equivalent plastic strain $\varepsilon_{eq,p}$, illustrating shear band development. (b) Comparison with experimental data from Bui et al. [14], where the green dots and black triangles denote the experimentally observed failure surface and free surface, respectively. The simulated failure surface is defined by a threshold $\varepsilon_{eq,p}>0.03$.
@@ 4.2 Application of JAX-MPM for 3D cases
To evaluate the effectiveness of JAX-MPM in modeling complex 3D geophysical dynamics, we consider three benchmark scenarios: (i) a 3D dam-break with embedded rigid obstacles (Section 4.2.1), (ii) 3D granular column collapse (Section 4.2.2), and (iii) 3D granular cylinder collapse (Section 4.2.3). The corresponding simulation configurations and computational costs are summarized in Table 1.
4.2.1 Simulation of 3D dam-break problem
As illustrated in Fig. 7, the computational domain for the 3D dam-break scenario is defined as $0,\,1.0]\times[0,\,0.5]\times[0,\,0.4]\,\mathrm{m}^{3}$. The initial dimensions of the water column are $H_{0}=0.2\,\mathrm{m}$ (height), $L_{0}=0.3\,\mathrm{m}$ (length), and $W_{0}=0.5\,\mathrm{m}$ (width). The fluid is modeled as a weakly compressible Newtonian fluid with a dynamic viscosity of $\mu=1.01\times 10^{-3}\,\mathrm{Pa\cdot s}$ and a numerical sound speed of $c_{n}=35\,\mathrm{m/s}$.
Two rigid square columns are placed in the domain as internal obstacles. Their side lengths are 0.04 m and 0.06 m, positioned at $(0.5,0.3,0.0)\,\mathrm{m}$ and $(0.6,0.1,0.0)\,\mathrm{m}$, respectively. A frictionless contact model employing a predictor–corrector scheme [8] is used to prevent fluid penetration. Since the rigid obstacles are stationary, contact enforcement reduces to nullifying the fluid’s normal velocity component at the interface:
@@ 4.2.1 Simulation of 3D dam-break problem
Figure 7: Initial configuration of 3D dam-break simulation with two rigid internal obstacles.
@@
Besides capturing physically realistic flow dynamics, JAX-MPM showcases excellent computational efficiency. The simulation presented in Fig. 8, featuring over 1.9 million particles and a $200\times 100\times 80$ grid. It runs on GPU with JIT compilation and completes 1000 time steps in approximately 8 seconds (see Table 1). This computational performance underscores the scalability and efficiency of the JAX-based implementation.
Figure 8: Snapshots of 3D dam-break simulation with rigid internal obstacles. The sequence from $t=0.12$ to $0.60$ s captures the fluid front impacting the columns, generating upward splashing, and subsequently forming recirculation zones and lateral flow separation downstream.
Table 1: Summary of simulation configurations and computational costs. The label “sim” refers to forward simulations, with runtimes reported for both single precision (float32) and double precision (float64), formatted as “float32 / float64”. The label “sim + grad” denotes inverse simulations involving both forward and gradient computations, performed exclusively in double precision (float64). “s/1000 steps” indicates the wall time (seconds) for every 1000 time steps.
Example Particle # Grid resolution Param # s/1000 steps $\Delta h$
Dam-break (Fig. 8) 1,920,000 $200\times 100\times 80$ N/A 8.2/35.1 (sim) $1/200$
@@
To evaluate the capability of JAX-MPM in simulating large-scale granular flow behaviors and capturing the influence of geometric variation, we consider the 3D granular column with varying initial aspect ratios. The objective is to investigate how the geometry, characterized by the aspect ratio $a=H_{0}/L_{0}$, affects the flow dynamics, runout distance, and failure mechanisms. Here, $H_{0}$ and $L_{0}$ denote the initial height and base length of the granular column, respectively. Simulations are carried out for a total of 200,000 time steps, corresponding to a physical duration of 2 s with a time step size of $\Delta t=10^{-5}\,\text{s}$. The material properties remain the same as in Section 4.1.2.
Four aspect ratios are investigated: $a=0.5,1.0,2.0,$ and $3.0$. As shown in Fig. 9, increasing the aspect ratio results in a larger portion of the column undergoing free fall, leading to reduced static basal zones and enhanced lateral spreading.
Moreover, the final runout distances $L_{f}$ are compared by defining the normalized runout as $d_{n}=(L_{f}-L_{0})/L_{0}$. The results show a clear increase in $d_{n}$ with aspect ratio: 2.05 ($a=0.5$), 3.97 ($a=1.0$), 7.76 ($a=2.0$), and 10.65 ($a=3.0$). The trend transitions from linear to sublinear scaling, consistent with experimental observations by Staron and Hinch [86], where $d_{n}\propto a$ for small $a$ and $d_{n}\propto a^{0.7}$ for larger values.
@@
To assess the computational performance, we benchmark JAX-MPM against the open-source C++ solver CB-Geo [59]. Five discretization resolutions are tested, with grid spacings ranging from $\Delta h=0.015$ to $\Delta h=0.004$, corresponding to particle counts from approximately $1.7\times 10^{4}$ to $9.0\times 10^{5}$. Wall-clock times are recorded for every 1000 simulation steps under the following hardware configurations: (1) CB-Geo on an AMD EPYC 7763 CPU with 128 threads, and (2) JAX-MPM on an NVIDIA A100 GPU. Note that CB-Geo employs a Mohr–Coulomb material model, as the built-in solver does not currently support Drucker–Prager plasticity. While Mohr–Coulomb models are generally more expensive due to their non-smooth yield surfaces, the computational complexity is comparable, and thus the comparison still provides meaningful insight into relative scalability.
[…]
Figure 9: Simulation results of 3D granular column collapse with varying aspect ratios. 2D deposit profiles are shown at $t=0.4\,\text{s},\,0.8\,\text{s},\,1.2\,\text{s},\,\text{and}\,\text{final deposit}$ with initial aspect ratios $a=H_{0}/L_{0}=0.5,1.0,2.0,3.0$. Higher aspect ratios lead to greater vertical collapse and longer runout distances due to enhanced free-fall and mass mobilization.
@@
We further assess the capability of JAX-MPM by simulating the collapse of a 3D granular cylinder with varying internal friction angles: $\varphi=20^{\circ},25^{\circ},35^{\circ},$ and $40^{\circ}$. Each simulation runs for 100,000 steps (1 s of physical time) with a time step size of $\Delta t=10^{-5}\,\text{s}$. The Drucker–Prager elastoplastic model is employed, with all other material parameters remaining identical to those detailed in Section 4.1.2.
Fig. 11 shows that increasing the friction angle produces steeper, more confined deposits and shorter runout, indicating greater resistance to shear deformation. These trends are consistent with theoretical and experimental studies reported in the literature [58, 56] and highlight the friction angle as a primary control on shear resistance, flow mobility, and final deposit geometry.
Computational performance discussion
The JAX-MPM simulation involves over 2.7 million particles on a $400\times 400\times 100$ computational grid (approximately 16 million grid nodes). On an NVIDIA A100 GPU, JAX-MPM completes 1000 simulation steps in approximately 22 seconds (in single precision) and 98 seconds (in double precision). For comparison, a recent high-performance Fortran-based MPM solver with dynamic load balancing (DLB) reported a runtime of 688 seconds for a similar scenario on 288 CPU cores, with a comparable material model and a grid resolution of $280\times 280\times 120$ (approximately 9.4 million grid nodes) [45]. Without DLB, the runtime increases significantly to roughly 13,803 seconds. These comparisons show that JAX-MPM achieves a $7\times$ speedup over the optimized Fortran+DLB implementation and over $140\times$ speedup relative to the baseline CPU version without DLB.
Together with the benchmarking results in Section 4.2, these comparisons highlight the computational efficiency and scalability of JAX-MPM. The framework’s performance is enabled by just-in-time (JIT) compilation, GPU parallelism, and kernel fusion (see Section 3.3.1), making it well suited for large-scale geomechanical simulations.
@@
In the following section, we further demonstrate the unique capability of JAX-MPM to perform inverse modeling through end-to-end differentiation, a task that is often cumbersome or intractable using conventional solvers.